I do not understand how to come up with the answer or an equation to help me get the answer for this problem. The word problem reads:
If 14 workers can complete a job in 2 days, how long will it take 4 workers to complete the same job? And assume that all workers work at the same rate.
I have tried to divided it multiply it and subtract but nothing makes sense to me. Please help!!!!!

Answer by stanbon(57377)

If 14 workers can complete a job in 2 days, how long will it take 4 workers to complete the same job? And assume that all workers work at the same rate.
---
-----------
# of Men and # of days are inversely related
d = k/m
---
Solve for "k" using "14 workers can complete a job in 2 days"
2 = k/14
k = 28
-----------
Then the equation relating men and days is
days = 28/men
-----
Problem: How long will it take 4 workers to complete the same job? And assume that all workers work at the same rate.
d = 28/4
# of days = 7
====================
Cheers,
Stan H.